OpenStudy (anonymous):

Find the inverse of: f(x)=lnx+4

OpenStudy (xapproachesinfinity):

to find the inverse first look up at this: - one-to-one property (Horizontal line test) - write the function as y=f(x) - switch y and x roles so that x=f(y) -then solve for y

OpenStudy (xapproachesinfinity):

he we have f(x)=lnx+4 is 1 to 1 so no problem here y=lnx+4 then x=lny+4 (switching roles) solve for y now

OpenStudy (anonymous):

\[ y= \ln x + 4\\ y-4 = \ln x\\ x=?? \]

OpenStudy (anonymous):

\[ x=e^{y-4} \] The inverse function is \[ y=e^{4-x} \]